# Preferential Attachment Trees with Vertex Death: Persistence of the Maximum Degree

### v1

@@ 4.3. Preliminary results
$=\bigg{(}\prod_{\ell=1}^{m}\mathbb{P}\!\left(D\geq v_{\ell}\right)\bigg{)}\mathbb{P}\!\left(\exists j\in\mathbb{N}_{0}:\sum_{\ell=1}^{m}\sum_{i=1}^{v_{\ell-1}}X(uv_{1}\ldots v_{\ell-1}i)+\sum_{i=1}^{K+j}X(uvi)\leq\sum_{i=K+1}^{K+j}X(Vi)\right),$
where the final step holds almost surely, since all random variables in the probability are independent of $\mathcal{F}_{T}$. As a result, by using Lemma 4.8 we arrive at
$\mathbb{E}_{\mathcal{S}}\big{[}{}$ $|\mathcal{D}^{(u)}|\,|\,\mathcal{F}_{T}\big{]}$ (4.85)
$\leq 1+\bigg{(}\prod_{i=K+1}^{\infty}\mathbb{E}\left[\mathrm{e}^{\lambda(X(i)-X(i)^{\prime})}\right]\bigg{)}\sum_{m=1}^{\infty}\sum_{v_{1}=1}^{\infty}\cdots\sum_{v_{m}=1}^{\infty}\prod_{\ell=1}^{m}\mathbb{P}\!\left(D\geq v_{\ell}\right)\mathbb{E}\left[\mathrm{e}^{-\lambda\sum_{i=1}^{v_{\ell}}X(i)}\right]$
$=1+\bigg{(}\prod_{i=K+1}^{\infty}\mathbb{E}\left[\mathrm{e}^{\lambda(X(i)-X(i)^{\prime})}\right]\bigg{)}\sum_{m=1}^{\infty}\mathbb{E}\left[\sum_{j=1}^{D}\mathrm{e}^{-\lambda\sum_{i=1}^{j}X(i)}\right]^{m}$
@@
$=1+\bigg{(}\prod_{i=K+1}^{\infty}\mathbb{E}\left[\mathrm{e}^{\lambda(X(i)-X(i)^{\prime})}\right]\bigg{)}\frac{\widehat{\mu}(\lambda)}{1-\widehat{\mu}(\lambda)},$
which is finite by the conditions in (PHC). As a result, since $A^{+}$ is finite $\mathbb{P}_{\mathcal{S}}$-almost surely,
[…]
Combining all of the above, we conclude that $\mathcal{P}_{1}$ is finite $\mathbb{P}_{\mathcal{S}}$-almost surely. ∎
We provide a proof of Proposition 4.14 for $m=2$ and later discuss the necessary adaptations when $m\geq 3$. As these are mainly a matter of notation and careful bookkeeping, we decided to provide the proof in the case $m=2$ rather than for general $m>1$ to aid the reader. We recall that in the case $m\geq 2$ we work under the additional assumption that the inter-birth times are exponentially distributed.
Proof of Proposition 4.14 for $m=2$.
We recall $K$ from (PHC) and define
$T:=\inf\{t\geq 0:|\{u\in\mathcal{U}_{\infty}:D^{(u)}=\infty,\mathcal{B}(uK)\leq t\}|\geq 2\}$ (4.87)
to be the stopping time at which the branching process contains at least two individuals that have produced $K$ many children and that satisfy that their ‘eventual’ offspring is infinite. Let $V_{1}$ and $V_{2}$ denote these two individuals, where we order them such that $\mathcal{B}(V_{1}K)<\mathcal{B}(V_{2}K)=T$. We observe that conditionally on survival such individuals exist and that $T$ is finite almost surely, since $\mathbb{P}\!\left(D=\infty\right)>0$. We now define
$\mathcal{A}(K):=\mathcal{U}_{K}^{<\infty}\cap\mathrm{BP}(T).$ (4.88)
We also define
$\mathcal{A}^{+}:=\bigg{(}\bigcup_{u\in\mathrm{BP}(T)}\{u1,\ldots,uI(u)\}\bigg{)}\setminus\mathrm{BP}(T),$ (4.89)
where we recall $I(u)$ from (4.72). In words, $\mathcal{A}^{+}$ consists of the individuals that are not yet born by time $T$ among the first $I(u)$ children of individuals $u\in\mathrm{BP}(T)$.
If $u\in\mathcal{P}_{2}$, this implies that there are two individuals $w_{1},w_{2}\in\mathrm{Pre}(u)$ such that $u$ catches up to either $w_{1}$ or to $w_{2}$. Indeed, if $u$ does not catch up to both, then $u$ can not catch up to all but at most one predecessor. Hence, for any $u\in\mathcal{P}_{2}$ we need to identify two predecessors and $u$ needs to catch up to either one of them. We let c.u.t. be short for ‘catches up to’ for ease of writing. As a result, we have
$\mathcal{P}_{2}\subseteq\mathrm{BP}(T){}$ $\cup\bigcup_{u\in\mathcal{A}^{+}}\{uv:v\in\mathcal{U}_{\infty},D^{(uv)}=\infty,uv\text{ c.u.t.\ }V_{1}\text{ or }V_{2}\}$ (4.90)
$\cup\!\!\!\!\!\!\!\!\!\!\bigcup_{\begin{subarray}{c}u\in\mathcal{A}(K)\setminus\{V_{1}\}\\ D^{(u)}=\infty\end{subarray}}\!\!\!\!\!\!\!\!\!\!\{uv:v\in\mathcal{U}_{\infty},v_{1}>K,D^{(uv)}=\infty,uv\text{ c.u.t.\ }u\text{ or }V_{2}\}$
$\cup\{V_{1}v:v\in\mathcal{U}_{\infty},v_{1}>K,\mathcal{B}(V_{1}v)>T,D^{(V_{1}v)}=\infty,V_{1}v\text{ c.u.t.\ }V_{1}\text{ or }V_{2}\}.$
Let us take some time to argue why this inclusion holds. First, we observe that we can partition $\mathcal{U}_{\infty}$ into all individuals born up to time $T$, i.e. $\mathrm{BP}(T)$, and all individuals born after time $T$. The three unions on the right-hand side of (4.90) consider all individuals born after time $T$. Furthermore, each set in one of the two big unions, as well as the final set on the right-hand side, identifies two predecessors to which an individual $uv$ or $V_{1}v$ has to catch up. We thus conclude that the right-hand side contains $\mathcal{P}_{2}$. To show that $\mathcal{P}_{2}$ is finite $\mathbb{P}_{\mathcal{S}}$-almost surely, it thus suffices to show that the right-hand side is finite $\mathbb{P}_{\mathcal{S}}$-almost surely.
To this end, we first observe that $\mathrm{BP}(T)$ is finite $\mathbb{P}_{\mathcal{S}}$-almost surely, by Lemma 4.7 and the fact that $T$ is finite $\mathbb{P}_{\mathcal{S}}$-almost surely. It then directly follows that $\mathcal{A}(K)\subseteq\mathrm{BP}(T)$ is also finite $\mathbb{P}_{\mathcal{S}}$-almost surely. Since each individual in $\mathrm{BP}(T)$ contributes only a finite number of elements to $\mathcal{A}^{+}$, it follows that $\mathcal{A}^{+}$ is finite $\mathbb{P}_{\mathcal{S}}$-almost surely as well. Furthermore, the sets $\mathrm{BP}(T)$, $\mathcal{A}^{+}$, and $\mathcal{A}(K)$ are all $\mathcal{F}_{T}$ measurable, where $\mathcal{F}_{T}$ is the $\sigma$-algebra generated by the process up to time $T$. We aim to prove that the right-hand side of (4.90) is finite conditionally on $\mathcal{F}_{T}$. As $\mathcal{A}^{+}$ and $\mathcal{A}(K)$ are $\mathcal{F}_{T}$ measurable and finite $\mathbb{P}_{\mathcal{S}}$ almost surely, we are hence left to argue that each of the sets in the unions is finite conditionally on $\mathcal{F}_{T}$. These follow from similar arguments, with minor differences only.
Let us start with the sets
$\{uv:v\in\mathcal{U}_{\infty},D^{(uv)}=\infty,uv\text{ c.u.t.\ }V_{1}\text{ or}V_{2}\},$ (4.91)
where $u$ is in $\mathcal{A}^{+}$. We prove that for any $u\in A^{+}$ and $j\in\{1,2\}$,
[…]
We obtain this by showing that the conditional expectation of the size of the set if finite. We write
$\mathbb{E}_{\mathcal{S}}{}$ [|$\Big{[}|\{uv:v\in\mathcal{U}_{\infty},D^{(uv)}=\infty,uv\text{ catches up to }V_{j}\}|\,\Big{|}\,\mathcal{F}_{T}\Big{]}$ (4.93)
$=\sum_{r=1}^{\infty}\sum_{v_{1}=1}^{\infty}\cdots\sum_{v_{r}=1}^{\infty}\mathbb{P}_{\mathcal{S}}\left(D^{(uv)}=\infty,uv\text{ c.u.t.\ }V_{j}\,\Big{|}\,\mathcal{F}_{T}\right).$
We define
$K_{j}:=\max\{k\in\mathbb{N}:\mathcal{B}(V_{j}k)\leq T\}\qquad\text{for }j\in\{1,2\},$ (4.94)
as the last child $V_{j}$ gave birth to before time $T$. Note that $K_{1}\geq K$ and $K_{2}=K$ $\mathbb{P}_{\mathcal{S}}$-almost surely by definition of $V_{1}$ and $V_{2}$. We then write the conditional probability as
$\mathbb{P}_{\mathcal{S}}\Bigg{(}{}$ $\exists n\in\mathbb{N}_{0}:\mathcal{B}(u)+\sum_{\ell=1}^{r}\sum_{i=1}^{v_{\ell}}X(uv_{1}\cdots v_{\ell-1}i)+\!\!\sum_{i=1}^{K_{j}+n}\!\!X(uvi)\leq\mathcal{B}(V_{j}K_{j})+\!\!\!\!\sum_{i=K_{j}+1}^{K_{j}+n}\!\!\!\!X(V_{j}i),$ (4.95)
$D^{(uv)}=\infty,D^{(uv_{1}\cdots v_{\ell-1})}\geq v_{\ell}\text{for all }\ell\in[r]\,\Bigg{|}\,\mathcal{F}_{T}\Bigg{)}.$
We now use that $u\in\mathcal{A}^{+}$, so that $u$ is not born by time $T$. At the same time, $V_{j}K_{j}$ is born by time $T$. That is, $\mathcal{B}(u)>T\geq\mathcal{B}(V_{j}K_{j})$. As a result, for $j=2$ we obtain the upper bound
$\mathbb{P}_{\mathcal{S}}\Bigg{(}\exists n\in\mathbb{N}_{0}:\sum_{\ell=1}^{r}\sum_{i=1}^{v_{\ell}}X(uv_{1}\cdots v_{\ell-1}i)+\!\!\sum_{i=1}^{K+n}\!\!X(uvi)\leq\!\!\!\!\sum_{i=K+1}^{K+n}\!\!\!\!X(V_{2}i)\,\Bigg{|}\,\mathcal{F}_{T}\Bigg{)}\mathbb{P}\!\left(D=\infty\right)\prod_{\ell=1}^{r}\mathbb{P}\!\left(D\geq v_{\ell}\right),$ (4.96)
where we also use that $K_{2}=K$ and that the offspring random variables $D^{(uv)}$ and $D^{(uv_{1}\cdots v_{\ell-1})}$ for $\ell\in[r]$ are independent of $\mathcal{F}_{T}$, as well as independent of the inter-birth random variables. Moreover, all inter-birth times in this upper bound are independent of $\mathcal{F}_{T}$ as well, since individual $V_{2}K$ is born exactly at time $T$ by definition. As a result, we can omit the conditioning on $\mathcal{F}_{T}$ $\mathbb{P}_{\mathcal{S}}$-almost surely to arrive at
$\mathbb{P}_{\mathcal{S}}\Bigg{(}\exists n\in\mathbb{N}_{0}:\sum_{\ell=1}^{r}\sum_{i=1}^{v_{\ell}}X(uv_{1}\cdots v_{\ell-1}i)+\!\!\sum_{i=1}^{K+n}\!\!X(uvi)\leq\!\!\!\!\sum_{i=K+1}^{K+n}\!\!\!\!X(V_{2}i)\Bigg{)}\mathbb{P}\!\left(D=\infty\right)\prod_{\ell=1}^{r}\mathbb{P}\!\left(D\geq v_{\ell}\right).$ (4.97)
When $j=1$, we instead have the upper bound
$\mathbb{P}_{\mathcal{S}}{}$ $\Bigg{(}\exists n\in\mathbb{N}_{0}:\sum_{\ell=1}^{r}\sum_{i=1}^{v_{\ell}}X(uv_{1}\cdots v_{\ell-1}i)+\!\!\sum_{i=1}^{K_{1}+n}\!\!X(uvi)\leq\mathcal{B}(V_{1}K_{1})-T+\!\!\!\!\sum_{i=K_{1}+1}^{K_{1}+n}\!\!\!\!X(V_{1}i)\,\Bigg{|}\,\mathcal{F}_{T}\Bigg{)}$ (4.98)
×$\times\mathbb{P}\!\left(D=\infty\right)\prod_{\ell=1}^{r}\mathbb{P}\!\left(D\geq v_{\ell}\right).$
Here, the inter-birth time $X(V_{1}(K_{1}+1))$ (as well as $T$ and $K_{1}$) depends on $\mathcal{F}_{T}$. However, by the memoryless property of the exponential distribution, it follows that $\mathcal{B}(V_{1}K_{1})+X(V_{1}(K_{1}+1))-T\sim X(V_{1}(K_{1}+1))^{\prime}$, where $X(V_{1}(K_{1}+1))^{\prime}$ is an i.i.d. copy of $X(V_{1}(K_{1}+1))$, which is independent of $\mathcal{F}_{T}$. As a result, we arrive at the equivalent upper bound as for $j=2$ $\mathbb{P}_{\mathcal{S}}$-almost surely, namely
$\mathbb{P}_{\mathcal{S}}\Bigg{(}\exists n\in\mathbb{N}_{0}:\sum_{\ell=1}^{r}\sum_{i=1}^{v_{\ell}}X(uv_{1}\cdots v_{\ell-1}i)+\!\!\sum_{i=1}^{K_{1}+n}\!\!X(uvi)\leq\!\!\!\!\sum_{i=K_{1}+1}^{K_{1}+n}\!\!\!\!X(V_{1}i)\,\Bigg{|}\,\mathcal{F}_{T}\Bigg{)}\mathbb{P}\!\left(D=\infty\right)\prod_{\ell=1}^{r}\mathbb{P}\!\left(D\geq v_{\ell}\right).$ (4.99)
Here, only the random variable $K_{1}$ depends on $\mathcal{F}_{T}$, all other random variables are independent of $\mathcal{F}_{T}$. Lemma 4.8 then yields for both $j=1$ and $j=2$ the upper bound
@@
$\prod_{i=K_{j}+1}^{\infty}\!\!\mathbb{E}\left[\mathrm{e}^{\lambda(X(i)^{\prime}-X(i))}\right]\prod_{i=1}^{K_{j}}\mathbb{E}\left[\mathrm{e}^{-\lambda X(i)}\right]\prod_{\ell=1}^{r}\Bigg{(}\mathbb{P}\!\left(D\geq v_{\ell}\right)\mathbb{E}\left[\exp\bigg{(}-\lambda\sum_{i=1}^{v_{\ell}}X(i)\bigg{)}\right]\Bigg{)},$ (4.100)
where we recall that $K_{2}=K$ is deterministic. However, we can almost surely obtain a deterministic upper bound by using that $K_{j}\geq K$ for both $j\in\{1,2\}$, which yields
@@
$\prod_{i=K+1}^{\infty}\!\!\mathbb{E}\left[\mathrm{e}^{\lambda(X(i)^{\prime}-X(i))}\right]\mathbb{E}\left[\exp\bigg{(}-\lambda\sum_{i=1}^{K}X(i)\bigg{)}\right]\prod_{\ell=1}^{r}\Bigg{(}\mathbb{P}\!\left(D\geq v_{\ell}\right)\mathbb{E}\left[\exp\bigg{(}-\lambda\sum_{i=1}^{v_{\ell}}X(i)\bigg{)}\right]\Bigg{)}.$ (4.101)
Using this in (4.93), we arrive at
$\mathbb{E}_{\mathcal{S}}{}$ [|$\Big{[}|\{uv:v\in\mathcal{U}_{\infty},D^{(uv)}=\infty,uv\text{ c.u.t.\ }V_{j}\}|\,\Big{|}\,\mathcal{F}_{T}\Big{]}$ (4.102)
[…]
@@
×$\times\sum_{r=1}^{\infty}\sum_{v_{1}=1}^{\infty}\cdots\sum_{v_{r}=1}^{\infty}\prod_{\ell=1}^{r}\Bigg{(}\mathbb{P}\!\left(D\geq v_{\ell}\right)\mathbb{E}\left[\exp\bigg{(}-\lambda\sum_{i=1}^{v_{\ell}}X(i)\bigg{)}\right]\Bigg{)}$
[…]

### v2

@@ 4.2. Preliminary results
$K_{j}\coloneq\sup\{k\in\mathbb{N}\colon\mathcal{B}(V_{j}k)\leq T\}\qquad\text{for }j\in\{1,2\}.$ (4.90)
Here, we note that $K_{2}=K$ by definition of $T$ and $V_{2}$, and $K_{1}\geq K$. We then write
$\mathbb{E}_{\mathcal{S}}{}$ $\Big[|\{V_{2}v\colon v\in\mathcal{U}_{\infty},v_{1}>K,D^{(V_{2}v)}=\infty,V_{2}v\text{ c.u.t.\ }V_{1}\}|\,\Big|\,\mathcal{F}_{T}\Big]$ (4.91)
$=\sum_{m=1}^{\infty}\sum_{v_{1}=K+1}^{\infty}\sum_{v_{2}=1}^{\infty}\cdots\sum_{v_{m}=1}^{\infty}\mathbb{P}_{\mathcal{S}}\left(D^{(V_{2}v)}=\infty,V_{2}v\text{ c.u.t.\ }V_{1}\,|\,\mathcal{F}_{T}\right).$
Since $K_{2}=K$ by definition of $T$ and $V_{1},V_{2}$, we can write each probability as
$\mathbb{P}_{\mathcal{S}}\bigg({}$ $\exists j\in\mathbb{N}_{0}\colon\mathcal{B}(V_{2}K)+\sum_{i=K}^{v_{1}-1}E^{(V_{2})}_{i}+\sum_{\ell=2}^{m}\sum_{i=0}^{v_{\ell}-1}E^{(V_{2}v_{1}\cdots v_{\ell-1})}_{i}+\!\!\sum_{i=0}^{K_{1}+j}\!E^{(V_{2}v)}_{i}\leq\mathcal{B}(V_{1}K_{1})+\!\!\sum_{i=K_{1}}^{K_{1}+j}\!E^{(V_{1})}_{i}$, (4.92)
$D^{(V_{2}v)}=\infty,D^{(V_{2}v_{1}\cdots v_{\ell-1})}\geq v_{\ell}\text{ for all }\ell\in\{2,\ldots,m\}\,\bigg|\,\mathcal{F}_{T}\bigg).$
Since $\mathcal{B}(V_{1}K_{1})\leq\mathcal{B}(V_{2}K)=T\leq\mathcal{B}(V_{1}(K_{1}+1))$ by the definition of $T$ and $K_{1},K_{2}$, and by omitting the event that $D^{(V_{2}v)}=\infty$, we can bound this probability from above by
$\mathbb{P}_{\mathcal{S}}\bigg({}$ $\exists j\in\mathbb{N}_{0}\colon\!\sum_{i=K}^{v_{1}-1}\!E^{(V_{2})}_{i}+\sum_{\ell=2}^{m}\sum_{i=0}^{v_{\ell}-1}E^{(V_{2}v_{1}\cdots v_{\ell-1})}_{i}+\!\!\!\sum_{i=1}^{K_{1}+j}\!E^{(V_{2}v)}_{i}\leq\mathcal{B}(V_{1}K_{1})-T+\!\sum_{i=K_{1}}^{K_{1}+j}\!E^{(V_{1})}_{i}$, (4.93)
$D^{(V_{2}v_{1}\cdots v_{\ell-1})}\geq v_{\ell}\text{ for all }\ell\in\{2,\ldots,m\}\,\bigg|\,\mathcal{F}_{T}\bigg).$
Only the random variables $V_{1},V_{2},K_{1},T$, and $E^{(V_{1})}_{K_{1}}$ depend on $\mathcal{F}_{T}$. All other random variables are independent of $\mathcal{F}_{T}$. By the memoryless property of the exponential distribution and the strong Markov property, we have
$\mathcal{B}(V_{1}K_{1})-T+E^{(V_{1})}_{K_{1}}\sim E_{K_{1}}^{\prime},$ (4.94)
where $E_{K_{1}}^{\prime}$ is an independent copy of $E^{(V_{1})}_{K_{1}}$, which is also independent of $\mathcal{F}_{T}$. The probability thus equals
$\mathbb{P}_{\mathcal{S}}\bigg({}$ $\exists j\in\mathbb{N}_{0}\colon\sum_{i=K}^{v_{1}-1}\!E^{(V_{2})}_{i}+\sum_{\ell=2}^{m}\sum_{i=0}^{v_{\ell}-1}E^{(V_{2}v_{1}\cdots v_{\ell-1})}_{i}+\!\!\!\sum_{i=1}^{K_{1}+j}\!E^{(V_{2}v)}_{i}\leq E_{K_{1}}^{\prime}+\!\sum_{i=K_{1}}^{K_{1}+j}\!E^{(V_{1})}_{i}$, (4.95)
$D^{(V_{2}v_{1}\cdots v_{\ell-1})}\geq v_{\ell}\text{ for all }\ell\in\{2,\ldots,m\}\,\bigg|\,\mathcal{F}_{T}\bigg).$
We now apply Lemma 4.3 (where we note that the condition in (4.8) is satisfied for any $\lambda>0$ by choosing $K$ large enough, see Remark 4.6), to obtain the upper bound
@@
$\frac{1}{\mathbb{P}\!\left(\mathcal{S}\right)}{}$ $\prod_{i=K_{1}}^{\infty}\mathbb{E}\left[\mathrm{e}^{\lambda(E_{i}^{\prime}-E_{i})}\right]\bigg(\prod_{i=0}^{K_{1}}\mathbb{E}\left[\mathrm{e}^{-\lambda E_{i}}\right]\bigg)\mathbb{P}\!\left(D\geq v_{1}\right)\mathbb{E}\left[\exp\bigg(-\lambda\sum_{i=K}^{v_{1}-1}E_{i}\bigg)\right]$ (4.96)
$\times\prod_{\ell=2}^{m}\mathbb{P}\!\left(D\geq v_{\ell}\right)\mathbb{E}\left[\exp\bigg(-\lambda\sum_{i=0}^{v_{\ell}-1}E_{i}\bigg)\right].$
Since $\mathbb{E}\left[\exp(\lambda(E_{i}^{\prime}-E_{i}))\right]\geq 1$ by Jensen’s inequality and $\mathbb{E}\left[\exp(-\lambda E_{i})\right]\leq 1$ for any $i\in\mathbb{N}$ and as $K_{1}\geq K$ by definition, we obtain an upper bound when replacing $K_{1}$ with $K$ in the first two products. Then, by the independence of the inter-birth times, we have
$\bigg(\prod_{i=0}^{K}\mathbb{E}\left[\mathrm{e}^{-\lambda E_{i}}\right]\bigg)\mathbb{P}\!\left(D\geq v_{1}\right)\mathbb{E}\left[\exp\bigg(-\lambda\!\!\!\sum_{i=K}^{v_{1}-1}\!\!\!E_{i}\bigg)\right]=\mathbb{P}\!\left(D\geq v_{1}\right)\mathbb{E}\left[\exp\bigg(-\lambda\sum_{i=0}^{v_{1}-1}E_{i}\bigg)\right].$ (4.97)
The right-hand side can thus be incorporated in the product on the second line of (4.96), so that this expression (with $K_{1}$ bounded by $K$) simplifies to
@@
$\frac{1}{\mathbb{P}\!\left(\mathcal{S}\right)}{}\prod_{i=K}^{\infty}\mathbb{E}\left[\mathrm{e}^{\lambda(E_{i}^{\prime}-E_{i})}\right]\prod_{\ell=1}^{m}\mathbb{P}\!\left(D\geq v_{\ell}\right)\mathbb{E}\left[\exp\bigg(-\lambda\sum_{i=0}^{v_{\ell}-1}E_{i}\bigg)\right].$ (4.98)
Using this upper bound in (4.91) and writing $c=1/\mathbb{P}\!\left(\mathcal{S}\right)$, we thus obtain
$\mathbb{E}_{\mathcal{S}}\Big[{}$ $|\{V_{2}v\colon v\in\mathcal{U}_{\infty},v_{1}>K,D^{(V_{2}v)}=\infty,V_{2}v\text{ c.u.t.\ }V_{1}\}|\,\Big|\,\mathcal{F}_{T}\Big]$ (4.99)
@@
$\leq{}$ $\sum_{m=1}^{\infty}\sum_{v_{1}=K+1}^{\infty}\sum_{v_{2}=1}^{\infty}\cdots\sum_{v_{m}=1}^{\infty}c\prod_{i=K}^{\infty}\mathbb{E}\left[\mathrm{e}^{\lambda(E_{i}^{\prime}-E_{i})}\right]\prod_{\ell=1}^{m}\mathbb{P}\!\left(D\geq v_{\ell}\right)\mathbb{E}\left[\exp\bigg(-\lambda\sum_{i=0}^{v_{\ell}-1}E_{i}\bigg)\right].$
Switching summation and products and extending the range of the variable $v_{1}$ to $v_{1}\in\mathbb{N}$ yields the upper bound
@@
$c\prod_{i=K}^{\infty}\mathbb{E}\left[\mathrm{e}^{\lambda(E_{i}^{\prime}-E_{i})}\right]\sum_{m=1}^{\infty}\Bigg(\sum_{v=1}^{\infty}\mathbb{P}\!\left(D\geq v\right)\mathbb{E}\left[\exp\bigg(-\lambda\sum_{i=0}^{v-1}E_{i}\bigg)\right]\Bigg)^{m}.$ (4.100)
By writing the inner sum as $\widehat{\mu}(\lambda)$ and using that, by Assumption (F-L), we can take $\lambda>0$ large enough so that $\widehat{\mu}(\lambda)<1$, as follows from (4.2), we conclude that the upper bound is finite. Using this in (4.69) combined with the earlier conclusion that the set $\{V_{2}v\colon v\in\mathcal{U}_{\infty},v_{1}>K,D^{(V_{2}v)}=\infty,V_{2}v\text{ c.u.t.\ }V_{1}\}$ is finite $\mathbb{P}_{\mathcal{S}}$-almost surely, we obtain
$\mathbb{P}_{\mathcal{S}}\left(|\{V_{2}v\colon v\in\mathcal{U}_{\infty},v_{1}>K,D^{(V_{2}v)}=\infty,V_{2}v\text{ c.u.t.\ }V_{1}\text{ or }V_{2}\,\Big|\,\mathcal{F}_{T}\right)=1\qquad\mathbb{P}_{\mathcal{S}}\text{-almost surely}.$ (4.101)
Sets in (4.70). This proof follows a similar approach as for the sets in (4.69). By conditioning on $\mathcal{F}_{T}$, we have
$\mathbb{E}_{\mathcal{S}}{}$ $\Big[|\{uv\colon v\in\mathcal{U}_{\infty},D^{(uv)}=\infty,uv\text{ c.u.t. }V_{j}\}|\,\Big|\,\mathcal{F}_{T}\Big]$ (4.102)
$=\sum_{m=1}^{\infty}\sum_{v_{1}=1}^{\infty}\cdots\sum_{v_{m}=1}^{\infty}\mathbb{P}_{\mathcal{S}}\left(D^{(uv)}=\infty,uv\text{ c.u.t. }V_{j}\,\Big|\,\mathcal{F}_{T}\right).$
Recalling $K_{j}$ from (4.90), observing that that $K_{1}\geq K$ and $K_{2}=K$ by definition of $T$ and $V_{1},V_{2}$, we can write each of the probabilities in (4.102) as
$\mathbb{P}_{\mathcal{S}}\bigg({}$ $\exists n\in\mathbb{N}_{0}\colon\mathcal{B}(u)+\sum_{\ell=1}^{m}\sum_{i=0}^{v_{\ell}-1}E^{(uv_{1}\cdots v_{\ell-1})}_{i}+\!\!\!\sum_{i=0}^{K_{j}+n}\!E^{(uv)}_{i}\leq\mathcal{B}(V_{j}K_{j})+\!\!\!\sum_{i=K_{j}}^{K_{j}+n}\!E^{(V_{j})}_{i}$, (4.103)
$D^{(uv)}=\infty,D^{(uv_{1}\cdots v_{\ell-1})}\geq v_{\ell}\text{ for all }\ell\in[m]\,\bigg|\,\mathcal{F}_{T}\bigg).$
Since $u\in\mathcal{A}_{+}$, we know that $\mathcal{B}(u)>T\geq\mathcal{B}(V_{j}K_{j})$, so that we obtain the upper bound
$\mathbb{P}_{\mathcal{S}}\bigg({}$ $\exists n\in\mathbb{N}_{0}\colon\sum_{\ell=1}^{m}\sum_{i=0}^{v_{\ell}-1}E^{(uv_{1}\cdots v_{\ell-1})}_{i}+\!\!\!\sum_{i=0}^{K_{j}+n}\!E^{(uv)}_{i}\leq\mathcal{B}(V_{j}K_{j})-T+\!\!\!\sum_{i=K_{j}}^{K_{j}+n}\!E^{(V_{j})}_{i}$, (4.104)
$D^{(uv)}=\infty,D^{(uv_{1}\cdots v_{\ell-1})}\geq v_{\ell}\text{ for all }\ell\in[m]\,\bigg|\,\mathcal{F}_{T}\bigg).$
Only the random variables $K_{j}$, $T$, $V_{j}$, and $E^{(V_{j})}_{K_{j}}$ depend on $\mathcal{F}_{T}$ (the latter only for $j=1$). Since the inter-birth times are exponentially distributed and by the strong Markov property, it follows that
$\mathcal{B}(V_{j}K_{j})+E^{(V_{j})}_{K_{j}}-T\sim E_{K_{j}}^{\prime}\qquad\text{for both }j\in\{1,2\},$ (4.105)
where $E_{K_{j}}^{\prime}$ is an i.i.d. copy of $E^{(V_{j})}_{K_{j}}$ that is independent of $\mathcal{F}_{T}$. In fact, for $j=2$, since $K_{2}=K$ we have $\mathcal{B}(V_{2}K_{2})+E^{(V_{2})}_{K_{2}}-T=E^{(V_{2})}_{K_{2}}$, which is independent of $\mathcal{F}_{T}$. For both $j=1$ and $j=2$, we thus obtain
$\mathbb{P}_{\mathcal{S}}\bigg({}$ $\exists n\in\mathbb{N}_{0}\colon\sum_{\ell=1}^{m}\sum_{i=0}^{v_{\ell}-1}E^{(uv_{1}\cdots v_{\ell-1})}_{i}+\!\!\!\sum_{i=0}^{K_{j}+n}\!E^{(uv)}_{i}\leq E_{K_{j}}^{\prime}+\!\!\!\sum_{i=K_{j}+1}^{K_{j}+n}\!E^{(V_{j})}_{i}$, (4.106)
$D^{(uv)}=\infty,D^{(uv_{1}\cdots v_{\ell-1})}\geq v_{\ell}\text{ for all }\ell\in[m]\,\bigg|\,\mathcal{F}_{T}\bigg).$
We can then apply Lemma 4.3 (where we note that the condition in (4.8) is satisfied for any $\lambda>0$ by choosing $K$ large enough, see Remark 4.6), to obtain the upper bound
$\frac{1}{\mathbb{P}\!\left(\mathcal{S}\right)}\prod_{i=K_{j}}^{\infty}\!\!\mathbb{E}\left[\mathrm{e}^{\lambda(E_{i}^{\prime}-E_{i})}\right]\prod_{i=0}^{K_{j}-1}\mathbb{E}\left[\mathrm{e}^{-\lambda E_{i}}\right]\prod_{\ell=1}^{m}\Bigg(\mathbb{P}\!\left(D\geq v_{\ell}\right)\mathbb{E}\left[\exp\bigg(-\lambda\sum_{i=0}^{v_{\ell}-1}E_{i}\bigg)\right]\Bigg).$ (4.107)
Since $\mathbb{E}\left[\exp(\lambda(E_{i}^{\prime}-E_{i}))\right]\geq 1$ by Jensen’s inequality and that $\mathbb{E}\left[\exp(-\lambda E_{i})\right]\leq 1$ for any $i\in\mathbb{N}$ and as $K_{j}\geq K$ for both $j\in\{1,2\}$, we can obtain an upper bound by replacing $K_{j}$ with $K$. Substituting this upper bound in (4.102) and writing $c=\mathbb{P}\!\left(\mathcal{S}\right)^{-1}$, we obtain
$\mathbb{E}_{\mathcal{S}}{}$ $\Big[|\{uv\colon v\in\mathcal{U}_{\infty},D^{(uv)}=\infty,uv\text{ c.u.t. }V_{j}\}|\,\Big|\,\mathcal{F}_{T}\Big]\leq\sum_{m=1}^{\infty}\sum_{v_{1}=1}^{\infty}\cdots\sum_{v_{m}=1}^{\infty}\prod_{i=K}^{\infty}\mathbb{E}\left[\mathrm{e}^{\lambda(E_{i}^{\prime}-E_{i})}\right]$ (4.108)
$c\times\prod_{i=0}^{K-1}\mathbb{E}\left[\mathrm{e}^{-\lambda E_{i}}\right]\prod_{\ell=1}^{m}\bigg(\mathbb{P}\!\left(D\geq v_{\ell}\right)\mathbb{E}\left[\exp\bigg(-\lambda\sum_{i=0}^{v_{\ell}-1}E_{i}\bigg)\right]\bigg).$
By switching summation and products, this equals
$c\prod_{i=K}^{\infty}\mathbb{E}\left[\mathrm{e}^{\lambda(E_{i}^{\prime}-E_{i})}\right]\prod_{i=0}^{K-1}\mathbb{E}\left[\mathrm{e}^{-\lambda E_{i}}\right]\sum_{m=1}^{\infty}\Bigg(\sum_{v=1}^{\infty}\mathbb{P}\!\left(D\geq v\right)\mathbb{E}\left[\exp\bigg(-\lambda\sum_{i=0}^{v-1}E_{i}\bigg)\right]\Bigg)^{m}.$ (4.109)
By writing the inner sum as $\widehat{\mu}(\lambda)$ and using that, by Assumption (F-L), we can take $\lambda>0$ large enough so that $\widehat{\mu}(\lambda)<1$, as follows from (4.2), we conclude that the upper bound is finite. Using this in (4.70), we thus have that
$\mathbb{P}_{\mathcal{S}}\left(|\{uv{}\colon v\in\mathcal{U}_{\infty},D^{(uv)}=\infty,uv\text{ c.u.t. }V_{1}\text{ or }V_{2}\}|<\infty\,\Big|\,\mathcal{F}_{T}\right)=1\qquad\mathbb{P}_{\mathcal{S}}\text{-almost surely}.$ (4.110)